OpenStudy (anonymous):

Establish: 1-sin^2 theta/1+cos^2 theta=cos theta

OpenStudy (anonymous):

Establish?

OpenStudy (anonymous):

does it help to know what \(1-\sin^2(\theta)=\cos^2(\theta)\)?

OpenStudy (agent0smith):

You'll have to use parentheses so we can tell exactly what it is...

OpenStudy (anonymous):



OpenStudy (anonymous):

it is better, but it is false

OpenStudy (anonymous):

I didn't solve it...supposedly the chapter is establishing trigonometric identities O.o

OpenStudy (anonymous):

it would be like saying \[\frac{5^2}{1+5^2}=5\] or \[\frac{x^2}{1+x^2}=x\]

OpenStudy (anonymous):

are you positive that is what it says? it is wrong for sure, but maybe it says something slightly different